OpenStudy (anonymous):

Nate invested three thousand eight hundred dollars in a certificate of deposit with a simple annual interest rate of four percent. How much interest will he earn after a year? one hundred fifty−two dollars nine hundred fifty dollars one thousand five hundred twenty dollars three thousand nine hundred fifty−two dollars

OpenStudy (nurali):

3800 X 4/100 X 1 = 152 $ ANSWER is A

OpenStudy (anonymous):

Karim borrowed one hundred twenty dollars from his older brother for the purchase of a skateboard. He promised to pay back the loan in 6 months at a simple annual interest rate of fifteen percent. How much will Karim owe his brother at that time? one hundred eight dollars one hundred twenty−nine dollars one hundred thirty−eight dollars two hundred twenty−eight dollars

OpenStudy (nurali):

120 X 15/100 X 1/2 = 9 120 + 9 = 129 $ ANSWER IS B)

OpenStudy (anonymous):

Russell deposited six hundred fifty dollars in a savings account that earns three percent simple annual interest. If he makes no other deposits or withdraws, what will the balance of Russell's account be after 6 years? one hundred seventeen dollars six hundred fifty−nine dollars seven hundred sixty−seven dollars one thousand one hundred seventy dollars

OpenStudy (anonymous):

thanls so much!!!!!!!!!!!!!!! please keep helping me!

OpenStudy (anonymous):

okkkkkkkk

OpenStudy (nurali):

50 X 3/100 X 6 = 11700/100 = 117 650 + 117 = 767 $ ANSWER is C

OpenStudy (anonymous):

almost done! Devante paid a security deposit of five hundred twenty−five dollars when he signed the lease for his apartment. The money was put into an account earning two percent simple annual interest. After renting for 8 years, Devante got his deposit back. What was the payout balance? four hundred forty−one dollars five hundred thirty−two dollars five hundred ninety−five dollars six hundred nine dollars

OpenStudy (nurali):

525 X 2/100 X 8 = 8400/100 = 84 525 +84 = 609 $ ANSWER is D

OpenStudy (anonymous):

Match the conditions given with the corresponding amount of interest. Question 1 – one thousand eight hundred dollars at two percent for ten years Question 2– one thousand dollars at six percent for four years Question 3– two thousand five hundred dollars at three percent for five years Question 4- nine hundred dollars at ten percent for three years OPTIONS- A) two hundred forty dollars B) three hundred seventy−five dollars C) three hundred sixty dollars D) two hundred seventy dollars

OpenStudy (anonymous):

@Nurali its the last question. lol

OpenStudy (nurali):

C = 360$ A = 240$ B= 375$ D= 270$

OpenStudy (anonymous):

THANK YOU SO MUCH!!!!!!!!!!!!!!!!!!!!!

OpenStudy (nurali):

Anytime.

OpenStudy (anonymous):

i have to do a 20 question quiz, can you help me with that today also ?

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

okay. here we go!!

OpenStudy (anonymous):

Four friends went scuba diving. According to the following statements, who spotted sea life at the fastest rate? Lance saw 4 reef sharks in 6 minutes. Jasmine saw 6 sea cucumbers in 8 minutes. Robert saw eleven jellyfish in fifteen minutes. Shilpa saw 8 pencil urchins in ten minutes.

OpenStudy (nurali):

Lance rate is 4/6 = 2/3 = 0.666 Jasmine rate is 6/8 = 3/4 = 0.75 Robert rate is 11/15 = 0.733333 Shilpa rate is 8/10 = 0.8 therefore answer is D

OpenStudy (anonymous):

A juice box contains 8 total ounces. If 4 ounces is fruit juice, which proportion could be used to determine what percentage of the juice box is actually fruit juice? eight over four equals x over one hundred one hundred over x equals four over eight x over eight equals four over one hundred x over one hundred equals four over eight

OpenStudy (nurali):

A juice box contains 8 total ounces. If 4 ounces is fruit juice, which proportion could be used to determine what percentage of the juice box is actually fruit juice? x over eight equals four over one hundred

OpenStudy (anonymous):

Curt's jazz ensemble drove to participate in a state music competition. They drove back without stopping, which decreased the travel time by ten percent, or thirty minutes. How many minutes was the drive to the competition? thirty-three minutes forty minutes two hundred seventy minutes three hundred minutes

OpenStudy (nurali):

30/x = 10/100 cross multiply 3000 = 10x x= 300

OpenStudy (anonymous):

Otto mixed spices to create his own Italian seasoning. He used ten percent garlic powder. Which of the following shows the percent equation and solution for how many tablespoons of garlic powder he uses to create twenty tablespoons of the seasoning? twenty equals zero point one times a; a equals two twenty equals zero point one times a; a equals two hundred a equals zero point one times twenty; a equals two hundred a equals zero point one times twenty; a equals two

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

a equals zero point one times twenty; a equals two

OpenStudy (anonymous):

Which of the following statements represents the highest amount? Income on a fifty thousand dollar salary after nineteen percent income tax Salary before a nineteen percent pay raise to fifty thousand dollars nineteen percent commission on fifty thousand dollars Salary before a nineteen percent pay cut to fifty thousand dollars

OpenStudy (nurali):

nineteen percent commission on fifty thousand dollars

OpenStudy (anonymous):

What percentage interest rate is a loan that charges one hundred dollars in simple annual interest over 5 years on a one thousand dollar balance? zero point two four percent two percent twenty percent twenty-four percent

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Nurali bless you for answering these questions.

OpenStudy (nurali):

Thanks @mosuechick0012

OpenStudy (anonymous):

Ok i took the fist quiz the ones you missed were Match the conditions given with the corresponding amount of interest. Question 1 – one thousand eight hundred dollars at two percent for ten years Question 2– one thousand dollars at six percent for four years Question 3– two thousand five hundred dollars at three percent for five years Question 4- nine hundred dollars at ten percent for three years OPTIONS- A) two hundred forty dollars B) three hundred seventy−five dollars C) three hundred sixty dollars D) two hundred seventy dollars

OpenStudy (anonymous):

how is it D the first question on the second quiz